# SWE-Dev: Evaluating and Training Autonomous Feature-Driven Software Development

### v2

@@ 3.3 Statistics
Table 2 summarizes the key statistics of SWE-Dev, which consists of 14,000 training and 500 test samples. The test set is manually curated and split into two difficulty levels: easy and hard (250 instances each). Each dataset instance comprises four components: (1) the task, specified by a PRD, with its token count reflecting instruction length; (2) the codebase, referring to the non-test portion of the repository, where we report the number of files and lines of code (LOC); (3) the ground truth (GT) code to be implemented, measured by its LOC and number of target functions; and (4) the test suite, evaluated via the number of test cases and total test LOC per sample. Figure 3 shows the distribution of training and test samples across six major PyPI application domains, demonstrating the diversity of software categories represented in the dataset. More statistics are in Appendix A.
4 Experiment
In this section, we empirically evaluate the effectiveness of various coding systems and training paradigms on SWE-Dev. We first compare the performance of single-LLM(§ 4.1) and MAS(§ 4.1.2) on the FDD tasks. Then, the effectiveness of different training approaches, including SFT (§ 4.2.1), RL (§ 4.2.2), and multi-agent training (§ 4.2.3) is discussed.
Setup. We employed the Pass@$k$ as an evaluation metric in SWE-Dev [20]. For inference code context, since SWE-Dev requires both the PRD and codebase as inputs. The codebases consist of many tokens (an average of 202K lines, see Table 2), exceeding typical LLM context window. Thus, in all the experiments below, we provide only the relevant code context—i.e., the specific files involved in the task—rather than the entire codebase.
4.1 Testing Results
This section presents the performance of 17 chatbot LLMs, 10 reasoning LLMs, and 10 multi-agent systems on SWE-Dev, under the single-LLM and multi-agent settings. Full details of the evaluated methods are provided in Appendix F.1.
Figure 4: Comparison of Pass@3 scores for 17 chatbot and 10 reasoning LLMs on SWE-Dev across Easy and Hard splits. SWE-Dev poses substantial challenges and effectively distinguishes model capabilities under both difficulty levels. See Appendix B for full results.
4.1.1 Single LLM Inference
SWE-Dev presents substantial challenges for current LLMs, revealing a clear gap between existing coding capabilities and real-world software engineering demands. Figure 4 reports Pass@3 performance of chatbot and reasoning LLMs on SWE-Dev. We observe that:
(1) LLMs perform better on the easy split than the hard split. (2) Performance generally scales with model size, especially for LLMs within the same family, aligning with our understanding of LLM capabilities. (3) Even the best-performing LLM (Claude-3.7-Sonnet[11]) achieves just over 20% on the Hard split. This still fall short of achieving strong performance, indicating that current models are not yet fully capable of handling tasks that approximate the complexity of real-world developing scenarios.
Reasoning models generally underperform their base counterparts, with Claude-3.7-Sonnet being a exception. While Claude with thinking outperforms its base variant, most reasoning models yield worse results. This suggests that current reasoning strategies do not consistently translate into gains for complex, repository-level generation tasks. We further explain this in Appendix C.2.
Figure 5: Comparison of benchmarks on various model sizes. SWE-Dev shows clear performance scaling with model size, while HumanEval[20] fails to distinguish between models. ComplexCodeEval[15] using CodeBLEU[33] exhibits high variance, making it less stable for evaluation.
Easy Hard
Method Pass@1 Calls Price(\$) Pass@1 Calls Price(\$)
Single 34.47 1.00 0.75 11.09 1.00 0.97
Reflexion [14] 39.77 2.12 0.83 13.32 2.18 1.35
Self Refine [34] 40.02 5.00 5.78 20.03 5.00 5.8
Self Consistency [35] 37.62 6.00 4.30 18.55 6.00 7.08
LLM Debate [36] 38.48 7.00 5.95 14.56 7.00 9.35
MAD [37] 31.50 7.00 2.48 15.31 7.00 3.40
Agentverse [38] 38.67 4.52 1.40 13.42 4.83 2.90
EvoMAC [18] 34.59 7.98 3.20 13.60 8.30 4.65
MetaGPT 29.56 9.69 2.20 9.25 10.37 4.95
MapCoder [39] 24.55 21.01 6.05 5.87 23.41 10.55
ChatDev [40] 35.13 26.61 3.53 11.70 30.87 6.10
SWE-Dev provides stable and discriminative evaluation of model capabilities.
Figure 5 compares the performance of Qwen2.5 [41] family on SWE-Dev, HumanEval [20], and ComplexCodeEval [15] across three runs. We use Pass@1 for SWE-Dev and HumanEval and ComplexCodeEval for CodeBLEU [33]. The lines represent the average performance, and the shaded regions show the variance. We observe that: (1) SWE-Dev yields low variance performance and consistent scaling with model size, demonstrating SWE-Dev’s stability and reliability in evaluating model capabilities. (2) In contrast, HumanEval—despite being stable—is too simple to differentiate models meaningfully. (3) Meanwhile, ComplexCodeEval shows high variance due to its reliance on similarity-based metrics, CodeBLEU, which limits its reliability for evaluating complex generation tasks.
4.1.2 Multi-Agent Inference
Table 5 compares the performance, call times and total costs of various MAS against the single-agent baseline driven by GPT-4o-mini. Details of MAS are given in AppendixF.1. Key observations are
MAS generally outperforms single-agent baselines on complex tasks. While the single-agent approach achieves only 11.09% Pass@1 on hard tasks, Self Refine[34] and EvoMAC[18] improve performance to 20.03% and 13.60%, respectively. These results highlight the advantage of multi-agent collaboration in solving complex, reasoning-intensive problems.
4.2 Training Results
In this section, we evaluate SWE-Dev’s support for different training methods, including SFT, RL. Additionally, we present preliminary results from our exploration of multi-agent training, offering an initial assessment of MAS-based learning. For detailed training setups, refer to the Appendix F.2.
4.2.1 Single LLM SFT
We conducted experiments on Qwen2.5-Intstruct models of various sizes (0.5B, 1.5B, 3B, and 7B) to assess the impact of SFT on performance in SWE-Dev. Experimental setting is in Appendix F.3.
Zero-shot SFT
Easy Hard Easy $\Delta$(%) Hard $\Delta$(%)
0.5B 6.39 1.00 12.12 +90 4.37 +337
1.5B 8.05 1.23 18.20 +126 7.64 +521
3B 15.93 5.27 27.53 +73 12.46 +136
7B 25.74 6.68 36.90 +43 18.89 +183
4.2.2 Single LLM RL
Pass@1 Pass@3
Method Easy Hard Easy Hard
vanilla 25.74 6.68 33.35 7.73
SFT 27.09 9.77 34.49 13.63
PPO (online RL) 28.30 12.25 32.69 14.33
DPO (offline RL) 25.89 10.36 31.32 14.66
SWE-Dev provides precise test cases enabling accurate rewards for coding tasks, supporting both online and offline RL. In this section, we explore the impact of RL on the Qwen2.5-7B-instruct using SWE-Dev. Considering the computational cost of RL, we limit our experiments in this section to 2k training samples. For full training setup, refer to the Appendix F.4.
Both online and offline RL improve performance, especially on hard tasks. Table 5 shows that both PPO [42] and DPO [43] significantly improve Pass@1 performance, especially on the Hard split. Furthermore, PPO outperforms SFT on the same training data. These findings highlight the advantages of RL training.
RL boosts one-shot success but not multi-sample gains. While RL fine-tuning yields improvements in Pass@1, it has minimal impact on Pass@3. Specifically, PPO achieves a Pass@1 of 28.30% on easy tasks, a noticeable increase from the base model’s 25.74%, but the Pass@3 remains lower than the SFT-training, even the original model’s performance. These results suggest that RL can be beneficial in refining Pass@1, particularly for more complex tasks, by increasing the model’s efficiency in generating correct answers in fewer attempts. However, this efficiency comes at the cost of reduced exploration. This aligns with findings from prior work[44]. Therefore, while RL improves performance, significant headroom remains, and more advanced methods or further training are needed to achieve improvements across tasks.
4.2.3 Multi-Agent Training
MAS has shown promising results on SWE-Dev, and we further investigate the training process of MAS on this dataset. As depicted in Fig. 7, the ground truth test case supervision in SWE-Dev enables EvoMAC [18] to improve its performance across multiple rounds of reasoning. This iterative refinement process motivates us to explore EvoMAC as the MAS for training in SWE-Dev. We apply rejection sampling to enhance agent performance via role-wise training.
Table 6: Comparison of multi-agent role-wise training, base MAS and single LLM’s performance on Qwen2.5-7B-Instruct. $\Delta$ indicates the relative improvement over the base MAS system. Partial fine-tuning of either agent also leads to consistent gains, demonstrating the effectiveness of role-specific supervision enabled by SWE-Dev.
Org Coder Easy $\Delta$(%) Hard $\Delta$(%)
Single - - 25.74 - 6.68 -
MAS base base 26.64 - 7.39 -
FT base 30.04 +12.76 12.36 +67.25
base FT 31.42 +17.94 11.49 +55.48
FT FT 31.65 +18.80 12.70 +71.85
5 Dataset Analysis
We analyze SWE-Dev’s task complexity, evaluation setup, and PRD quality to demonstrate its uniqueness and reliability.
(a) Complexity analysis
(b) Metric Comparison
(c) PRD Quality Analysis
Analysis of Task Difficulty and Call Tree Characteristics. We analyze how task difficulty in SWE-Dev correlates with call tree complexity. As introduced in § 3.1, a call tree reflects the dynamic function invocation structure for this task, where nodes represent functions and edges denote their call relationships. We use two metrics: depth, indicating the maximum call nesting, and node count, representing the total number of distinct functions involved in the task. Fig. 8 shows that GPT-4o’s performance declines as depth and node count increase, revealing a strong correlation between structural complexity and task difficulty. This suggests that deeper and broader call structures introduce more functional requirements and interdependencies, making tasks more challenging.
Evaluation Method Precision. SWE-Dev uses execution-based evaluation with test cases, enabling precise performance signals. We compare metrics: Exact Match (EM) [19], Exact Sequence (ES) [19], CodeBLEU [33], and Pass@3, using Qwen2.5 models and GPT-4o. As Fig 8 shows, Pass@3 best reflects capability scaling, separating models by size and quality. In contrast, EM, ES, and CodeBLEU show minimal variance, failing to distinguish models. This demonstrates that SWE-Dev’s test-case-based evaluation provides a more robust and realistic signal of model performance, better reflecting the functional correctness required in real-world software development.
PRD Quality. SWE-Dev includes a PRD for each task to simulate realistic developer-facing requirements, which are primarily derived from the original docstrings found within the repository source code. While many functions in open-source code include docstrings, we found that these are often incomplete—lacking clear descriptions of behavior, parameters, or edge cases. To improve instruction clarity without fabricating content, we lightly refine existing docstrings using GPT-4o, grounded in the related file and surrounding context. To evaluate instruction quality, we conducted a human assessment on 100 sampled tasks. Two experienced engineers rated the original and refined PRDs along Actionability, Completeness, and Clarity (Appendix C.1 includes human instruction). As shown in Fig. 8, refined PRDs consistently scored higher across all dimensions. This supports SWE-Dev’s goal of providing realistic, well-scoped requirements for reliable model evaluation.
6 Conclusion
In this work, we introduced SWE-Dev, the first dataset for evaluating and training autonomous coding systems on feature-driven development task. SWE-Dev consists of 14,000 training and 500 test instances, each uniquely equipped with runnable environments and developer-authored executable unit tests, which provides essential execution-based feedback for advanced training paradigms like SFT, RL, and multi-agent learning. Our experiments show FDD is profoundly challenging for current autonomous coding systems. We also validate that training on SWE-Dev can yield encouraging performance gains. These findings validate SWE-Dev as a critical platform for identifying current limitations and fostering breakthroughs in AI-driven software development. We hope the release of SWE-Dev spurs innovation in long-context reasoning, agent orchestration, and execution-aware training towards genuinely autonomous software engineering.
Limitation SWE-Dev currently focuses on Python and basic training strategies, leaving open future directions in multilingual support and more advanced training approaches.
Appendix A Dataset
A.1 Dataset Information
@@
PRD Description of the project development requirements for this task.
@@ A.3 Dataset Diversity
Package Category Diversity. To analyze the functional diversity of the dataset, we classify packages into high-level categories based on their primary domain (e.g., web development, data science, utilities). The classification is performed using GPT-4o-mini with the prompt provided below (see Figure 23). The resulting distribution confirms that SWE-Dev spans a broad spectrum of software domains.
@@
Appendix B Inference Results
To assess the capabilities and limitations of current LLMs on realistic feature-driven development (FDD) tasks, we conduct comprehensive inference-time evaluations on SWE-Dev. We study both single-agent and multi-agent systems, measuring their performance under consistent execution-based evaluation.
B.1 Single-Agent LLM Performance
We evaluate 27 state-of-the-art LLMs, including general-purpose chatbot models (e.g., GPT-4o, Claude 3.7) and reasoning models, shown in Table 9. Models are assessed using Pass@1 and Pass@3 on SWE-Dev’s test set. To contextualize benchmark difficulty, we also compare results on HumanEval [20] and ComplexCodeEval [15], using Pass@3 and CodeBLEU respectively. Our findings show that SWE-Dev poses significantly greater challenges than existing benchmarks, with leading models achieving under 25% Pass@3 on hard tasks (Table 8).
@@ B.1 Single-Agent LLM Performance
Model SWE-Dev pass@1 SWE-Dev pass@3 ComplexCodeEval CodeBLEU HumanEval pass@1 HumanEval pass@3
@@
Chatbot Qwen2.5-1.5B-Instruct 8.05% 1.23% 10.76% 2.22% 29.72% 57.20% 69.76%
Qwen2.5-3B-Instruct 15.93% 5.27% 21.99% 7.47% 12.27% 62.68% 75.00%
Qwen2.5-7B-Instruct 25.74% 6.68% 33.35% 7.73% 20.00% 82.68% 87.13%
Llama-3.1-8B-Instruct 26.43% 7.94% 33.01% 10.24% 20.18% 68.20% 77.87%
Qwen3-8B 34.04% 12.09% 39.26% 13.33% 17.47% 86.34% 89.09%
Qwen2.5-Coder-14B-Instruct 39.51% 14.82% 52.49% 18.44% 35.52% 90.48% 92.93%
Qwen2.5-14B-Instruct 38.08% 13.16% 46.32% 15.89% 19.90% 82.56% 87.87%
DeepSeek-Coder-V2-Lite-Instruct 21.53% 8.19% 29.68% 11.33% 26.63% 80.98% 85.18%
Qwen3-30B-A3B 35.84% 12.76% 39.45% 15.20% 14.84% 89.27% 90.30%
Phi-4 21.99% 5.57% 27.89% 8.56% 33.85% 86.46% 90.01%
Qwen2.5-32B-Instruct 43.64% 10.15% 51.24% 11.69% 19.76% 88.90% 92.44%
Qwen2.5-72B-Instruct 49.01% 10.62% 57.20% 12.33% 22.15% 83.66% 86.46%
Llama3.3-70B-Instrcut 33.84% 12.85% 39.57% 14.95% 21.29% 84.51% 88.54%
Deepseek-V3 41.95% 16.22% 56.79% 21.62% 28.32% 90.36% 92.92%
GPT-4o 54.37% 19.13% 68.70% 21.91% 33.38% 88.41% 92.93%
GPT-4o-mini 34.47% 11.09% 41.94% 13.84% 25.00% 85.97% 89.00%
Claude-3.7-Sonnet 53.09% 19.74% 56.35% 24.25% 29.63% 93.66% 95.36%
Reasoning Claude-3.7-Sonnet-thinking 49.47% 22.51% 56.58% 29.28% 29.80% 91.22% 97.62%
Deepseek-R1-distill-Qwen2.5-7B 6.30% 1.29% 10.30% 1.95% 21.05% 86.10% 93.29%
Qwen3-8B-thinking 19.47% 6.36% 25.91% 9.22% 20.98% 89.63% 91.89%
Qwen3-30B-A3B-thinking 23.63% 8.30% 31.00% 11.60% 25.00% 93.04% 99.57%
Deepseek-R1-distill-Qwen2.5-32B 24.25% 9.79% 40.53% 19.04% 27.98% 95.17% 97.87%
DeepSeek-R1-distill-Llama-70B 32.73% 8.19% 45.72% 11.33% 25.95% 96.95% 98.53%
Deepseek-R1-671B 28.55% 12.84% 37.62% 17.72% 34.47% 98.65% 100%
QwQ-32B-Preview 4.50% 0.70% 8.90% 1.22% 24.78% 82.31% 97.01%
grok-3-beta 53.63% 18.97% 59.08% 22.26% 27.96% 87.15% 89.99%
o1 36.36% 11.09% 43.77% 14.27% 33.63% 97.43% 98.78%
o3 51.21% 21.86% 59.05% 28.98% 26.53% 98.04% 98.78%
@@
B.2 Multi-Agent System Performance
We evaluate 10 multi-agent systems (MAS), including both general-purpose MAS (e.g., AgentVerse, LLM Debate) and code-specific designs (e.g., EvoMAC, MapCoder). As detailed in Table 5, we compare each MAS against a single-agent baseline on execution success (Pass@1), total API call count, and cost-efficiency. Results show that while MAS can outperform single agents on complex tasks, simple strategies (e.g., Self-Refine) often strike a better balance between performance and resource usage than workflow-heavy systems like ChatDev. We visualize the MAS in the following Figure 13.
@@ B.2 Multi-Agent System Performance
Appendix C Analysis
C.1 Analysis of PRD Quality

### v3

@@ C.1 Dataset Statistics
Table 5 summarizes the key statistics of SWE-Dev, which consists of 14,000 training and 500 test samples. The test set is manually curated and split into two difficulty levels: easy and hard (250 instances each). Each dataset instance comprises four components: (1) the task, specified by a PRD, with its token count reflecting instruction length; (2) the codebase, referring to the non-test portion of the repository, where we report the number of files and lines of code (LOC); (3) the ground truth (GT) code to be implemented, measured by its LOC and number of target functions; and (4) the test suite, evaluated via the number of test cases and total test LOC per sample. Figure 6 shows the distribution of training and test samples across six major PyPI application domains, demonstrating the diversity of software categories represented in the dataset. More statistics are in Appendix C.
C.2 Dataset Information
@@
PRD A natural language document describing the project requirements, including the specific features to be implemented.
@@ C.4 Dataset Diversity
Package Category Diversity. To analyze the functional diversity of the dataset, we classify packages into high-level categories based on their primary domain (e.g., web development, data science, utilities). The classification is performed using GPT-4o with the prompt provided below (see Figure 25). The resulting distribution confirms that SWE-Dev spans a broad spectrum of software domains.
@@
Appendix D Inference Results
To assess the capabilities and limitations of current LLMs on realistic feature-driven software development tasks, we conduct comprehensive inference-time evaluations on SWE-Dev. We study both single-agent and multi-agent systems, measuring their performance under consistent execution-based evaluation.
D.1 Single-Agent LLM Performance
We evaluate 27 state-of-the-art LLMs, including general-purpose chatbot models (e.g., GPT-4o, Claude 3.7) and reasoning models, shown in Table 8. Models are assessed using Pass@1 and Pass@3 on SWE-Dev’s test set. To contextualize benchmark difficulty, we also compare results on HumanEval [humaneval] and ComplexCodeEval [complexcodeeval], using Pass@3 and CodeBLEU respectively. Our findings show that SWE-Dev poses significantly greater challenges than existing benchmarks, with leading models achieving under 30% Pass@3 on hard tasks (Table 7).
@@ D.1 Single-Agent LLM Performance
Abbreviated Name Model SWE-Dev pass@1 SWE-Dev pass@3 ComplexCodeEval CodeBLEU HumanEval pass@1 HumanEval pass@3
@@
Chatbot Qwen2.5-1.5B Qwen2.5-1.5B-Instruct 8.05% 1.23% 10.76% 2.22% 29.72% 57.20% 69.76%
Qwen2.5-3B Qwen2.5-3B-Instruct 15.93% 5.27% 21.99% 7.47% 12.27% 62.68% 75.00%
Qwen2.5-7B Qwen2.5-7B-Instruct 25.74% 6.68% 33.35% 7.73% 20.00% 82.68% 87.13%
Llama3.1-8B Llama-3.1-8B-Instruct 26.43% 7.94% 33.01% 10.24% 20.18% 68.20% 77.87%
Qwen3-8B Qwen3-8B 34.04% 12.09% 39.26% 13.33% 17.47% 86.34% 89.09%
Qwen2.5Coder-14B Qwen2.5-Coder-14B-Instruct 39.51% 14.82% 52.49% 18.44% 35.52% 90.48% 92.93%
Qwen2.5-14B Qwen2.5-14B-Instruct 38.08% 13.16% 46.32% 15.89% 19.90% 82.56% 87.87%
DeepseekCoder-V2-Lite DeepSeek-Coder-V2-Lite-Instruct 21.53% 8.19% 29.68% 11.33% 26.63% 80.98% 85.18%
Qwen3-30B Qwen3-30B-A3B 35.84% 12.76% 39.45% 15.20% 14.84% 89.27% 90.30%
Phi-4 Phi-4 21.99% 5.57% 27.89% 8.56% 33.85% 86.46% 90.01%
Qwen2.5-32B Qwen2.5-32B-Instruct 43.64% 10.15% 51.24% 11.69% 19.76% 88.90% 92.44%
Qwen2.5-72B Qwen2.5-72B-Instruct 49.01% 10.62% 57.20% 12.33% 22.15% 83.66% 86.46%
Llama3.3-70B Llama3.3-70B-Instruct 33.84% 12.85% 39.57% 14.95% 21.29% 84.51% 88.54%
Deepseek-V3 Deepseek-V3 41.95% 16.22% 56.79% 21.62% 28.32% 90.36% 92.92%
GPT-4o GPT-4o 54.37% 19.13% 68.70% 21.91% 33.38% 88.41% 92.93%
gpt-4o-mini GPT-4o-mini 34.47% 11.09% 41.94% 13.84% 25.00% 85.97% 89.00%
Claude-3.7-Sonnet Claude-3.7-Sonnet 53.09% 19.74% 56.35% 24.25% 29.63% 93.66% 95.36%
Reasoning Claude-3.7-Sonnet(T) Claude-3.7-Sonnet-thinking 49.47% 22.51% 56.58% 29.28% 29.80% 91.22% 97.62%
Deepseek-R1(Qwen2.5-7B) Deepseek-R1-distill-Qwen2.5-7B 6.30% 1.29% 10.30% 1.95% 21.05% 86.10% 93.29%
Qwen3-8B(T) Qwen3-8B-thinking 19.47% 6.36% 25.91% 9.22% 20.98% 89.63% 91.89%
Qwen3-30B(T) Qwen3-30B-A3B-thinking 23.63% 8.30% 31.00% 11.60% 25.00% 93.04% 99.57%
Deepseek-R1(Qwen2.5-32B) Deepseek-R1-distill-Qwen2.5-32B 24.25% 9.79% 40.53% 19.04% 27.98% 95.17% 97.87%
Deepseek-R1(Llama-70B) DeepSeek-R1-distill-Llama-70B 32.73% 8.19% 45.72% 11.33% 25.95% 96.95% 98.53%
Deepseek-R1 Deepseek-R1-671B 28.55% 12.84% 37.62% 17.72% 34.47% 98.65% 100%
QwQ-32B QwQ-32B-Preview 4.50% 0.70% 8.90% 1.22% 24.78% 82.31% 97.01%
Grok-3 grok-3-beta 53.63% 18.97% 59.08% 22.26% 27.96% 87.15% 89.99%
o1 o1 36.36% 11.09% 43.77% 14.27% 33.63% 97.43% 98.78%
o3 o3 51.21% 21.86% 59.05% 28.98% 26.53% 98.04% 98.78%
@@
D.2 Multi-Agent System Performance
We evaluate 10 multi-agent systems (MAS), including both general-purpose MAS (e.g., AgentVerse, LLM Debate) and code-specific designs (e.g., EvoMAC, MapCoder). As detailed in Table D.2, we compare each MAS against a single-agent baseline on execution success (Pass@1), total API call count, and cost-efficiency. Results show that while MAS can outperform single agents on complex tasks, simple strategies (e.g., Self-Refine) often strike a better balance between performance and resource usage than workflow-heavy systems like ChatDev. We visualize the MAS in the following Figure 11.
@@ D.2 Multi-Agent System Performance
Easy Hard
Method Pass@1 Calls Price(\$) Pass@1 Calls Price(\$)
Single 34.47 1.00 0.75 11.09 1.00 0.97
Reflexion 39.77 2.12 0.83 13.32 2.18 1.35
Self Refine 40.02 5.00 5.78 20.03 5.00 5.8
Self Consistency 37.62 6.00 4.30 18.55 6.00 7.08
LLM Debate 38.48 7.00 5.95 14.56 7.00 9.35
MAD 31.50 7.00 2.48 15.31 7.00 3.40
Agentverse 38.67 4.52 1.40 13.42 4.83 2.90
EvoMAC 34.59 7.98 3.20 13.60 8.30 4.65
MetaGPT 29.56 9.69 2.20 9.25 10.37 4.95
MapCoder 24.55 21.01 6.05 5.87 23.41 10.55
ChatDev 35.13 26.61 3.53 11.70 30.87 6.10
D.3 Agent LLM Results
Table 10: Performance of tool-augmented LLM agents on SWE-Dev using the OpenHands framework. We report Pass@1 (%) on the Easy and Hard splits. Abbreviated names match those used in the main figures.
Abbreviated Name Model Name Easy Hard
Claude4.1-Opus(T) Claude-4.1-Opus-Thinking 86.30 56.44
GPT-5 GPT-5 87.56 54.21
Claude4.5(T) Claude 4.5 Thinking 84.83 53.55
KimiK2(T) Kimi-K2-Thinking 83.89 50.86
KimiK2 Kimi-K2-Instruct 81.82 51.05
GLM4.5 GLM-4.5 71.83 39.00
Qwen3-235B(T) qwen3-235b-a22b-thinking-2507 43.37 19.78
Qwen3-30B(T) qwen3-30b-a3b-thinking-2507 29.68 6.11
Agent Evaluation Setup.
All agent-based experiments are conducted using the OpenHands [wang2024openhands] framework. We use a unified configuration for all agent LLMs: temperature =0.6=0.6, maximum agent steps =100=100, execution timeout =300=300 seconds per task, and maximum context length =64=64K tokens. Performance is measured by executing the original SWE-Dev unit tests in the provided environments and reporting Pass@1 on the Easy and Hard splits.
Appendix E Analysis
E.1 Evaluation Metric Validness
SWE-Dev provides stable and discriminative evaluation of model capabilities. Figure 12 compares the performance of Qwen2.5 [qwen2.5] family on SWE-Dev, HumanEval [humaneval], and ComplexCodeEval [complexcodeeval] across three runs. We use Pass@1 for SWE-Dev and HumanEval and ComplexCodeEval for CodeBLEU [codebleu]. The lines represent the average performance, and the shaded regions show the variance. We observe that: (1) SWE-Dev yields low variance performance and consistent scaling with model size, demonstrating SWE-Dev’s stability and reliability in evaluating model capabilities. (2) In contrast, HumanEval—despite being stable—is too simple to differentiate models meaningfully. (3) Meanwhile, ComplexCodeEval shows high variance due to its reliance on similarity-based metrics, CodeBLEU, which limits its reliability for evaluating complex generation tasks.
Figure 12: Comparison of benchmarks on various model sizes. SWE-Dev shows clear performance scaling with model size, while HumanEval [humaneval] fails to distinguish between models. ComplexCodeEval [complexcodeeval] using CodeBLEU [codebleu] exhibits high variance, making it less stable for evaluation.
E.2 Analysis of PRD Quality